# Galactic echoes

### v1

@@ Acknowledgements
We are grateful to N. Bahcall, U. Banik, J. Goodman and M. Medvedev for helpful feedback at various stages of this project. R.C. is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference #DIS-2022-568580) and the Japan Society for the Promotion of Science (JSPS) Research Fellowship. C.H. is supported by the John N. Bahcall Fellowship Fund and the Sivian Fund at the Institute for Advanced Study.
@@ Appendix A Net kick by passing satellite in the impulsive regime
We consider the perturbation by a satellite galaxy and/or a dark-matter subhalo with mass $M_{\rm p}$ passing through the Solar neighborhood on a straight orbit with constant velocity $v_{\rm p}$ at an incident angle $\phi_{\rm p}$ between the velocity vector and the $z$-axis. We assume that the perturber is a Plummer sphere with core radius $a$. The potential of the perturber at the position of stars with height $z$ in the Solar neighborhood at time $t$ is
$\Phi_{\rm p}(z,t)=-\frac{{\rm G}M_{\rm p}}{\sqrt{(v_{\rm p}t\sin\phi_{\rm p})^{2}+(z-v_{\rm p}t\cos\phi_{\rm p})^{2}+a^{2}}}.$ (31)
@@
$\Delta v_{z}(z)$ $=-\int_{-\infty}^{\infty}\mathrm{d}t\frac{\mathrm{d}\Phi_{\rm p}}{\mathrm{d}z}$
$=-\int_{-\infty}^{\infty}\mathrm{d}t\frac{{\rm G}M_{\rm p}(z-v_{\rm p}t\cos\phi_{\rm p})}{[(v_{\rm p}t\sin\phi_{\rm p})^{2}+(z-v_{\rm p}t\cos\phi_{\rm p})^{2}+a^{2}]^{3/2}}.$ (32)
When the perturber passes at a glancing angle $\phi_{\rm p}\approx\pi/2$, the kick reduces to
$\Delta v_{z}(z)=-\int_{-\infty}^{\infty}\mathrm{d}t\frac{{\rm G}M_{\rm p}z}{[(v_{\rm p}t)^{2}+z^{2}+a^{2}]^{3/2}}=-\frac{{\rm G}M_{\rm p}}{v_{\rm p}}\frac{2z}{z^{2}+a^{2}},$ (33)
which is an odd function of $z$, thus producing a quadrupole perturbation in phase space.
@@ Appendix B Nonlinear saturation of echo
Here we explore the behavior of the echo in the fully nonlinear (i.e., non-perturbative) regime. As in §2.2 we ignore both the self-gravity of the fluctuations and any small-scale diffusion. In addition, we consider a homogeneous system, i.e., a system with no steady-state gravitational potential (so this calculation is very similar to the corresponding plasma calculation in §IV of O’Neil & Gould 1968). The advantage of the homogeneous assumption is that we do not have to use angle-action variables.
@@
First, the density shows multiple echo-like structures just after the second impulse. Since these features are absent at small $\psi_{2}$, they are most likely higher-order echoes. While one could in principle derive explicit solutions for these echoes by expanding (37) in powers of $\psi_{1}$ and $\psi_{2}$, this becomes increasingly tedious at higher orders. However, the timing of these higher-order echoes can be inferred as follows. Each power of $\psi_{i}$ is associated with sines and cosines with argument $\phi_{i}\equiv k_{i}[x-v(t-t_{i})]$. The terms of order $l$ have a factor $\psi_{1}^{m}\psi_{2}^{l-m}$ where $m=1,...,l-1$ (the terms with $m=0$ and $m=l$ arise from a single hammer and therefore do not produce an echo). These terms involve $\sin^{m}\phi_{1}\sin^{l-m}\phi_{2}$ and analogous cosine terms and so produce terms like $\sin(j_{1}\phi_{1})\sin(j_{2}\phi_{2})$ where $0\leq j_{1}\leq m,0\leq j_{2}\leq l-m$ and analogous cosine terms. Expanding these products of sines and cosines, we get oscillatory terms with argument $j_{1}\phi_{1}\pm j_{2}\phi_{2}$. These can produce echoes at $t=t^{(j_{1},j_{2})}_{\mathrm{echo}}$ where
@@
With the upper sign, the second term is always negative, so the echo appears before $t_{2}$ and thus is unphysical [c.f. equation (23))]. Therefore echoes occur only at
@@
Another nonlinear feature found at large $\psi_{2}$ is the saturation of the second-order echo. Precisely, as $\psi_{2}$ is increased, the echo around $t=40(\sigma k)^{-1}$ initially grows in amplitude but then saturates and decays—a behavior also observed in our (inhomogeneous) simulations (Fig. 4). Interestingly, at even larger $\psi_{2}$, the echo grows again but the density perturbation flips sign. This behavior results from the higher-order terms with $j_{1}=j_{2}$, which produce echoes with the same timing as the second-order echo. At each order $l$, the dominant contribution comes from terms proportional to $\psi_{1}\psi_{2}^{l-1}$, since these include a term proportional to $(t_{2}-t_{1})^{l-1}$ and $t_{2}-t_{1}$ is a large parameter. Taking the $l$th derivative of (37)—once with respect to $\psi_{1}$ and $l{-}1$ times with respect to $\psi_{2}$–and retaining only terms proportional to $(t_{2}-t_{1})^{l-1}$, we find that

### v2

@@ Acknowledgements
We are grateful to N. Bahcall, U. Banik, J. Goodman and M. Medvedev for helpful feedback at various stages of this project. RC is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference #DIS-2022-568580) and the Japan Society for the Promotion of Science (JSPS) Research Fellowship, grant No. 25KJ0049. CH is supported by the John N. Bahcall Fellowship Fund and the Sivian Fund at the Institute for Advanced Study. MWK was supported in part by NSF CAREER Award No. 1944972.
@@
Computations were performed on the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto.
@@ Appendix A Net kick by passing satellite in the impulsive regime
Figure 6: Amplitude of phase spirals as a function of time. The second perturbation at $t=t_{2}$ is modeled either as impulsive (solid) or time-dependent (dotted). The upper panel shows the fiducial case with $a=0.3\,{\rm kpc}$ and $M_{\rm p}=1.2\times 10^{8}\,{\rm M}_{\odot}$, while the lower panel adopts a more extended and massive perturber with $a=1.5\,{\rm kpc}$ and $M_{\rm p}=10^{9}\,{\rm M}_{\odot}$. For larger $a$, the perturbation becomes less impulsive and hence the time-dependent effect becomes more pronounced.
In this appendix, we present our model for the net vertical kick induced by the second perturber on a low-inclination orbit using the impulse approximation and discuss its validity. For an earlier study on satellite perturbations in a 1D stellar disc model, see Weinberg (1991).
We consider the perturbation by a satellite galaxy and/or a dark-matter subhalo with mass $M_{\rm p}$ passing through the Solar neighborhood at time $t=t_{\rm p}$. We assume that the perturber is on a straight orbit with constant velocity $v_{\rm p}$ and incident angle $\vartheta_{\rm p}$ between the velocity vector and the $z$-axis. We further assume that the perturber is a Plummer sphere with core radius $a$. The potential of the perturber at the position of stars with height $z$ in the Solar neighborhood at time $\tau=t-t_{\rm p}$ from the impact time is
$\Phi_{\rm p}(z,\tau)=-\frac{{\rm G}M_{\rm p}}{\sqrt{(v_{\rm p}\tau\sin\vartheta_{\rm p})^{2}+(z-v_{\rm p}\tau\cos\vartheta_{\rm p})^{2}+a^{2}}}.$ (31)
@@
$\Delta v(z)=-\int_{-\infty}^{\infty}\mathrm{d}\tau\frac{\mathrm{d}\Phi_{\rm p}}{\mathrm{d}z}$
$=-\int_{-\infty}^{\infty}\mathrm{d}\tau\frac{{\rm G}M_{\rm p}(z-v_{\rm p}\tau\cos\vartheta_{\rm p})}{[(v_{\rm p}\tau\sin\vartheta_{\rm p})^{2}+(z-v_{\rm p}\tau\cos\vartheta_{\rm p})^{2}+a^{2}]^{3/2}}.$ (32)
When the perturber passes at a glancing angle $\vartheta_{\rm p}\approx\pi/2$, the kick reduces to
$\Delta v(z)=-\int_{-\infty}^{\infty}\mathrm{d}\tau\frac{{\rm G}M_{\rm p}z}{[(v_{\rm p}\tau)^{2}+z^{2}+a^{2}]^{3/2}}=-\frac{{\rm G}M_{\rm p}}{v_{\rm p}}\frac{2z}{z^{2}+a^{2}},$ (33)
which is an odd function of $z$, thus producing a quadrupole perturbation in phase space. The kick peaks at $z=\pm a$ with amplitude $\Delta V_{2}={\rm G}M_{\rm p}/(av_{\rm p})$.
To test the validity of the impulse approximation, we compare simulations based on equation (33) with those that evolve under the fully time-dependent perturbation (31). Throughout, we keep the first perturber impulsive and only investigate the time-dependent effect of the second. The upper panel of Fig. 6 shows the amplitude of the phase spirals for impulsive (solid) and time-dependent (dotted) perturbations using the default parameters (§3.1): for the first perturber, $\Delta V_{1}=6\,{\rm kpc}\,{\rm Gyr}^{-1}$, and for the second, $a=0.3\,{\rm kpc}$, $M_{\rm p}=1.2\times 10^{8}\,{\rm M}_{\odot}$, $\vartheta_{\rm p}=\pi/2$, and $v_{\rm p}=300\,{\rm kpc}\,{\rm Gyr}^{-1}$, which gives $\Delta V_{2}\simeq 6\,{\rm kpc}\,{\rm Gyr}^{-1}$. The time-dependent perturbation produces nearly the same result as the impulsive case since the characteristic time of the perturbation is $a/v_{\rm p}=1\,{\rm Myr}$, which is much smaller than the stars’ vertical oscillation period, $T\simeq 70\,{\rm Myr}$, justifying the impulse approximation. The amplitude $A_{2}$ is slightly reduced, however, because stars rotate in phase space by a few times $(a/v_{\rm p})(2\pi/T)\sim 5^{\circ}$ during the passage of the perturber, so the perturbing force partially averages out.
In our fiducial model, we assumed a rather compact second perturber ($a=0.3\,{\rm kpc}$, $M_{\rm p}=1.2\times 10^{8}\,{\rm M}_{\odot}$) intended to represent a dark subhalo passing in the vicinity of the Solar neighborhood. For a Sagittarius-like perturber, both its size and mass may be larger. The lower panel of Fig. 6 shows results with $a=1.5\,{\rm kpc}$ and $M_{\rm p}=10^{9}\,{\rm M}_{\odot}$, with other parameters unchanged. [^4] For reference, the present-day Sagittarius dwarf galaxy has mass $M_{\rm p}=4\times 10^{8}\,{\rm M}_{\odot}$ and core radius of roughly $4^{\circ}$ on the sky, or $a\simeq 1.7\,{\rm kpc}$ assuming a distance of $25\,{\rm kpc}$ (Vasiliev & Belokurov, 2020). Despite the higher mass, the resulting perturbation $A_{2}$ is weaker than in the fiducial case (upper panel), since the kick is now strongest at $z=\pm 1.5\,{\rm kpc}$, well beyond the disc’s scale height ($h=0.2\,{\rm kpc}$). The time-dependent perturbation still produces a sharp rise in amplitude, as $a/v_{\rm p}=5\,{\rm Myr}$ remains short compared to $T$, but it is suppressed significantly relative to the impulsive case because stars now advance in phase by a few times $(a/v_{\rm p})(2\pi/T)\sim 26^{\circ}$ while the perturber passes through.
@@ Appendix B Nonlinear saturation of echo
This appendix explores the behavior of the echo in the fully nonlinear (i.e., non-perturbative) regime. As in §2.2 we ignore both the self-gravity of the fluctuations and any small-scale diffusion. Furthermore, we consider a homogeneous system, i.e., a system with no steady-state gravitational potential (so this calculation is very similar to the corresponding plasma calculation in §IV of O’Neil & Gould 1968). The advantage of the homogeneous assumption is that we do not have to use angle-action variables.
@@
First, the density shows multiple echo-like structures just after the second impulse. Since these features are absent at small $\psi_{2}$, they are most likely higher-order echoes. While one could in principle derive explicit solutions for these echoes by expanding (37) in powers of $\psi_{1}$ and $\psi_{2}$, this becomes increasingly tedious at higher orders. However, the timing of these higher-order echoes can be inferred as follows. Each power of $\psi_{i}$ is associated with sines and cosines with argument $\phi_{i}\equiv k_{i}[x-v(t-t_{i})]$. The terms of order $l$ have a factor $\psi_{1}^{m}\psi_{2}^{l-m}$ where m=1,,l1m=1,...,l-1 (the terms with $m=0$ and $m=l$ arise from a single hammer and therefore do not produce an echo). These terms involve $\sin^{m}\phi_{1}\sin^{l-m}\phi_{2}$ and analogous cosine terms and so produce terms like $\sin(j_{1}\phi_{1})\sin(j_{2}\phi_{2})$ where $0\leq j_{1}\leq m,0\leq j_{2}\leq l-m$ and analogous cosine terms. Expanding these products of sines and cosines, we get oscillatory terms with argument $j_{1}\phi_{1}\pm j_{2}\phi_{2}$. These can produce echoes at $t=t^{(j_{1},j_{2})}_{\mathrm{echo}}$ where
@@
With the upper sign, the second term is always negative, so the echo appears before $t_{2}$ and thus is unphysical [c.f. equation (23)]. Therefore echoes occur only at
@@
Another nonlinear feature found at large $\psi_{2}$ is the saturation of the second-order echo. Precisely, as $\psi_{2}$ is increased, the echo around $t=40(\sigma k)^{-1}$ initially grows in amplitude but then saturates and decays—a behavior also observed in our (inhomogeneous) simulations (Fig. 4). Interestingly, at even larger $\psi_{2}$, the echo grows again but the density perturbation flips sign. This behavior results from the higher-order terms with $j_{1}=j_{2}$, which produce echoes with the same timing as the second-order echo. At each order $l$, the dominant contribution comes from terms proportional to $\psi_{1}\psi_{2}^{l-1}$, since these include a term proportional to $(t_{2}-t_{1})^{l-1}$ and $t_{2}-t_{1}$ is a large parameter. Taking the $l$th derivative of (37)—once with respect to $\psi_{1}$ and l1l{-}1 times with respect to $\psi_{2}$–and retaining only terms proportional to $(t_{2}-t_{1})^{l-1}$, we find that